# A finite element-based physics-informed operator learning framework for spatiotemporal partial differential equations on arbitrary domains

### v2

@@ 4.1 Influence of training samples
This subsection examines the effect of the training samples on the prediction accuracy over time. The analysis includes not only the combination of Gaussian, Fourier, and constant fields but also the combinations of two out of the three temperature generators. To ensure that the constant fields accounted for only of the total training samples, we determined the proportion of samples from each generator. For the dataset of the combination of Fourier and constant fields, we prepared Fourier-type samples and constant-type samples. Similarly, for the dataset of the combination of Gaussian and constant fields, we prepared Gaussian-type samples and constant-type samples. To observe the impact of sample size on prediction, a new case was added with training samples, each with the same percentage as the original one. For this study, the NNs were trained for epochs with [s]. We also kept this condition for the other studies in the following subsections as well. The results are shown in Figs. 17 and 18 for the homogeneous and heterogeneous thermal conductivities, respectively. Notably, the combination of Fourier and constant resulted in errors about twice as large as in the other three cases, including those obtained from samples. This suggests that the accuracy of the prediction increases as more of the possible temperature field patterns are covered by training samples. The comparison of the temperature fields at [s] with the initial temperature field suggests that the fluctuation in the predictions concerning the reference solution by FEM is mitigated by enhancing training samples. When we employed the homogeneous thermal conductivity, it is clear that the error sometimes decreased from the previous time steps. This can happen in the present framework since the input temperature fields at the subsequent time steps after the first step are likely to be similar to some of the training samples, resulting in a better interpolation accuracy in the network prediction than the previous step. We also give a detailed discussion on this point in the next subsection about influence of time step size. For the heterogeneous conductivity case, the prediction accuracy with the combination of Fourier and constant was low compared to the other three types of training samples. This is the same trend as for the homogeneous case. The contours on the right of the figure show that the NNs failed to correctly predict the temperature evolution, particularly in the low conductivity regions when only the Fourier and constant generators are used; see Fig. 18 (a), which also means that the frequency in a variety of training samples greatly affects the prediction. Overall, the employment of diverse patterns for the training data sets was shown to be effective in improving the predictive performance of the present framework.
@@ 4.5 Influence of network architecture
The present framework consists of separate NNs for corresponding nodal temperatures at the next time step. On this matter, other network architecture options could be conceived, such as a fully connected one that returns the whole output field from a single NN. To quantitatively ensure the superiority of the present NN architecture, we evaluated the three types of network architectures, including the original one used for the main study [61]. The created network architectures are compared in Fig. 22. Compared to the original architecture, the elementwise-connected architecture takes nodal temperatures from adjacent elements as input for the output nodal temperature. For the fully connected architecture, nodal temperatures are provided as inputs and the entire nodal temperature field for the next time step is returned as output. In this study, 4 layers with 99 neurons in each layer, which is the number of nodes evaluated in the training, were selected. It is noted that the nodal number starts with 2 as we apply hard constraints for the Dirichlet boundary conditions and, therefore, eliminate those nodes from the training targets, as explained in Section 3.2. The results shown in Fig. 23 indicate that the two network architectures newly introduced here led to worse accuracy compared to the original architecture by a factor of 6. Separating NNs for different outputs while considering the entire problem domain for input is the most effective way to learn the mapping between the input and output physical fields in FOL. In the future, it would be interesting to investigate whether increasing the size of the element groups in Fig. 22 (b) can enhance predictive accuracy, which is in the end identical to that of Fig. 22 (a). Reducing the number of input dimensions to each NN shown in Fig. 22 (a) and (b) leads to the reduction of parameters in NNs in the present framework, resulting in less training cost. This improvement would play a role when one wants to apply FOL to a model with more nodes than what is considered in this work. Furthermore, it would also be beneficial to even reduce the number of neurons or layers in each network architecture, leading to the prevention of overfitting in NNs.
@@ 4.6 Influence of mesh size
Since the mesh size affects the prediction accuracy and the subsequent training cost due to the change in the number of required network parameters, it is also pivotal to gain insights on this point. For that, two different element sizes, one is the same one as in the main study () and the other is twice finer than the original one (), on the squared domain were considered to perform the comparison. Due to the increase in the training cost for finer mesh, we performed this study on the JAX platform which has the equivalent architecture to the SciANN-based code for faster training. The fully-connected network architecture with the number of neurons in each layer set to the number of nodes apart from the Dirichlet boundary nodes was employed for simplicity. The batch size was set to 60 for 3000 samples and 100 for 5000 samples to be consistent in terms of the batch size ratio to the total number of samples. The obtained average relative L2 errors along with the final are shown in Figs. 24 and 25 for the homogeneous and heterogeneous thermal conductivities, respectively. In Fig. 24, the magnitudes of the average relative L2 error norm over time did not exhibit major differences between and . One can also confirm that by enhancing training samples in the case of one can achieve a decent improvement in prediction accuracy. On the other hand, in Fig. 25 the error blew up after a few time steps when 3000 training samples were utilized for training. This was significantly mitigated by enhancing the training samples from 3000 to 5000, part of which is through adding higher frequencies to the Fourier series sample generator. This indicates that in the heterogeneous case, the quality of training samples critically affects the prediction accuracy in the finer mesh as shown in the case of , which is different from the observation in Fig. 18 where one does not see such an extreme error evolution. When it comes to the training cost, one has to pay the price for increasing the number of nodes in the finite element model as well as increasing the number of training samples. As summarized in Table 3, the homogeneous case with required 1.442 times more training time than the same setup with , whereas the heterogeneous case with experienced 2.664 times more training time to achieve as reasonable prediction accuracy as the same problem and network setup with due to the need for the enhancement of training samples. It is noted that the training was done within 30 minutes for all four cases using JAX. In application scenarios, one has to therefore adapt the variety of training samples to achieve reasonable prediction, especially when considering heterogeneous domains. Further exploring the integration of this method with data-driven auto-encoders presents intriguing possibilities to address potential constraints associated with increased mesh densities (i.e. higher resolutions), as highlighted in [66, 69, 70].
| Normalized training time | Required number of training samples | Relative L2 error | |
| Homogeneous, | 1.003 | 3000 | |
| Homogeneous, | 1.442 | 3000 | |
| Heterogeneous, | 1 | 3000 | |
| Heterogeneous, | 2.664 | 5000 |
@@ 4.7 Capability of handling arbitrary domains
shown in Figs. 27 and 28, the overall solution trend obtained by the proposed FOL framework agreed with the solution by FEM with the maximum absolute error around 0.1. The error was concentrated in the upper right and left part of the domain at and [s], which is also explained by the steep change in the temperature evolution due to the presence of heterogeneity. Furthermore, the error magnitudes did not differ very much from the results with the square domain as in Figs. 12, 13, and 14. The results of Fig. 28 are particularly noteworthy. This proves that FOL works for a problem setup with complex initial temperature, heterogeneity, and irregular domain discritized by unstructured mesh without cumbersome modification to the framework.
@@ 4.8 Computational cost and advantages of proposed framework
The main goal of this work is to establish a surrogate model for conventional numerical solvers. In this context, the prediction by the networks should be faster than the solution by numerical analysis. To quantitatively evaluate the speed of obtaining solutions by the FOL framework, the runtimes of the network inference with the same network architecture and finite element calculation were measured by performing the same task. The measurement was performed on the same CPU platform and environment to ensure the fairness of the measurement. We assumed ten-time steps in solving the heat equation with FEM. Therefore, the same network evaluation was performed ten times as well. As a result, the network prediction time was 10.8 times faster than FEM for the same setup. This result suggests that the network has the potential to be used as a surrogate for classic numerical solvers.
Although a faster inference than solving with a classical solver can be achieved with the proposed framework, one has to train the NNs for a relatively long time. For example, the training time for 1000 epochs on the problem setup shown in Figs. 3 and 6 took approximately 6 hours on a single GPU node of NVIDIA GeForce RTX 2080 12GB. However, the training time will be much faster using JAX (high-performance machine learning framework), which has superior features for deep learning, such as just-in-time compilation and vectorization. In addition, training NNs in FOL is usually a one-time investment. Once trained, users can use it for any input and can obtain the solution much quicker than numerical solvers, even for a model that requires many nodes to solve accurately.
@@ 5 Conclusion
This study has presented a novel physics-informed operator learning framework based on the finite element discretization scheme. After training with various temperature fields, including those generated by Gaussian distribution and Fourier series, as well as constant temperature fields, the network can accurately predict dynamic temperature evolutions for any arbitrary temperature input within the assumed temperature range. This is achieved with a relative L2 error below 0.1 in most cases, without the need for retraining under fixed boundary conditions and domain. The applicability of the method to heterogeneous heat conductivity and irregular domains is also confirmed. Additionally, the suggested network design can achieve over ten times the speed of the corresponding FEM solver on the same platform. It is important to note that the training is conducted entirely without ground truth data obtained from numerical simulations, making the framework a completely unsupervised learning approach. Furthermore, the training efficiency is enhanced compared to other operator learning approaches that rely on time-consuming automatic differentiation. This is because the current framework uses FE-based discretization for space and backward difference approximation for time. To summarize, this work explores the development of deep learning-based surrogates for dynamic physical phenomena without the need for supervised learning.
On the other hand, although the proposed framework offers useful features, there are still some limitations that could be addressed in future work. Firstly, heat conductivity can also be a target of training in addition to the temperature field. This makes the framework flexible for various micro morphologies with phase-field modeling in mind. It may be possible to improve accuracy by implementing a higher-order temporal discretization scheme.
One could also think about different network architectures that take multiple temperature fields as input.
Lastly, this study focused on transient heat conduction to showcase the performance of the framework. The present framework could be extended to other types of spatiotemporal PDEs, such as the convection-diffusion equation, the Allen-Cahn equation, or the Cahn-Hilliard equation, as the framework is developed with the aim of a generic deep learning framework for spatiotemporal dynamics phenomena described by PDEs.
@@ Acknowledgements:
Author Statement:
Y.Y.: Methodology, Software, Formal analysis, Investigation, Writing - Original Draft, A.H.: Conceptualization, Methodology, Writing - Original Draft, M.M.: Resources, Writing - Review & Editing, A.V.: Writing - Review & Editing, M.A.: Funding Acquisition, Writing -Review & Editing, T.B.: Writing - Review & Editing, S.R.: Funding Acquisition, Writing - Review & Editing, Sh.R.: Conceptualization, Methodology, Software, Supervision, Writing - Original Draft.

### v3

@@ 4.1 Influence of training samples
This subsection examines the effect of the training samples on the prediction accuracy over time. The analysis includes not only the combination of Gaussian, Fourier, and constant fields but also the combinations of two out of the three temperature generators. To ensure that the constant fields account for only of the total training samples, we determined the proportion of samples from each generator. For the dataset of the combination of Fourier and constant fields, we prepared Fourier-type samples and constant-type samples. Similarly, for the dataset of the combination of Gaussian and constant fields, we prepared Gaussian-type samples and constant-type samples. To observe the impact of sample size on prediction, a new case was added with training samples, each with the same percentage as the original one. For this study, the NNs were trained for epochs with [s]. We also kept this condition for the other studies in the following subsections as well. The results are shown in Figs. 18 and 19 for the homogeneous and heterogeneous thermal conductivities, respectively. Notably, the combination of Fourier and constant resulted in errors about twice as large as in the other three cases, including those obtained from samples. This suggests that the accuracy of the prediction increases as more of the possible temperature field patterns are covered by training samples. The comparison of the temperature fields at [s] with the initial temperature field suggests that the fluctuation in the predictions concerning the reference solution by FEM is mitigated by enhancing training samples. When we employed the homogeneous thermal conductivity, it was clear that the error sometimes decreased from the previous time steps. This can happen in the present framework since the input temperature fields at the subsequent time steps after the first step are likely to be similar to some of the training samples, resulting in a better interpolation accuracy in the network prediction than the previous step. We also give a detailed discussion on this point in the next subsection about the influence of time step size. For the heterogeneous conductivity case, the prediction accuracy with the combination of Fourier and constant was low compared to the other three types of training samples. This is the same trend as for the homogeneous case. The contours on the right of the figure show that the NNs failed to correctly predict the temperature evolution, particularly in the low conductivity regions when only the Fourier and constant generators are used; see Fig. 19 (a), which also means that the frequency in a variety of training samples greatly affects the prediction. Overall, the employment of diverse patterns for the training data sets was shown to be effective in improving the predictive performance of the present framework.
@@ 4.5 Influence of network architecture
The present framework consists of separate NNs for corresponding nodal temperatures at the next time step. On this matter, other network architecture options could be conceived, such as a fully connected one that returns the whole output field from a single NN. To quantitatively ensure the superiority of the present NN architecture, we evaluated the three types of network architectures, including the original one used for the main study [69]. The created network architectures are compared in Fig. 23. Compared to the original architecture, the elementwise-connected architecture takes nodal temperatures from adjacent elements as input for the output nodal temperature. For the fully connected architecture, nodal temperatures are provided as inputs and the entire nodal temperature field for the next time step is returned as output. In this study, 4 layers with 170 neurons in each layer were selected to ensure that the number of trainable parameters is comparable with each other. As a result, the fully connected architecture has 121,139 trainable parameters, whereas the separated and elementwise-connected architectures have 110,979 trainable parameters. It is noted that the nodal number starts with 2 as we apply hard constraints for the Dirichlet boundary conditions and, therefore, eliminate those nodes from the training targets, as explained in Section 3.2. The results shown in Fig. 24 indicate that the two network architectures newly introduced here led to worse accuracy compared to the original architecture approximately by a factor of 5. Separating NNs for different outputs while considering the entire problem domain for input is the most effective way to learn the mapping between the input and output physical fields in FOL. In terms of the training cost in the above cases, on the other hand, the fully connected architecture spent 6 hours 23 minutes 54 seconds, whereas the separated architecture took 7 hours 29 minutes 30 seconds and the elementwise connected architecture 7 hours 40 minutes 32 seconds when they were trained with SciANN for 1000 epochs on a single GPU node of NVIDIA GeForce RTX 2080 12GB, indicating that the fully connected one can be an efficient option that can still serve as a surrogate with sufficiently high-accuracy prediction. This is also supported by the comparison of the inference costs, where the fully connected architecture was shown to be 3.103 times faster than the separated network architecture in the measurement. Regarding the improvement of the separated and elementwise-
connected architectures, one interesting direction in the future would be to investigate whether increasing the size of the element groups in Fig. 23 (b) can enhance predictive accuracy, which is in the end identical to that of Fig. 23 (a). Reducing the number of input dimensions to each NN shown in Fig. 23 (a) and (b) leads to the reduction of parameters in NNs in the present framework, resulting in less training cost. This improvement would play a role when one wants to apply FOL to a model with more nodes than what is considered in this work. Furthermore, it would also be beneficial to even reduce the number of neurons or layers in each network architecture, leading to the prevention of overfitting in NNs. Nevertheless, for large-scale problems in which a large number of degrees of freedom need to be taken into account, one would have more benefits with the fully connected architecture, especially in combination with learning in latent space by employing techniques to condense information such as autoencoder.
@@ 4.6 Influence of mesh size
Since the mesh size affects the prediction accuracy and the subsequent training cost, it is also pivotal to gain insights on this point. For that, three different element sizes, one is the same one as in the main study () and the others are finer than the original one (), on the squared domain were considered to perform the comparison. Due to the increase in the training cost for finer mesh, we performed this study on the JAX platform which has the equivalent architecture to the SciANN-based code for faster training. The fully connected network architecture with the number of neurons in each layer set to 170 and 4 layers was employed for the study. The batch size was set to 60 for 3000 samples and 100 for 5000 samples to be consistent in terms of the batch size ratio to the total number of samples. The obtained average relative L2 errors along with the temperature distribution at s for the initial temperature field are shown in Figs. 25 and 26 for the homogeneous and heterogeneous thermal conductivities, respectively. In Fig. 25, the magnitudes of the average relative L2 error norm over time exhibited noticeable differences between , , and . This could be due to the reduced number of trainable parameters in each nodal evaluation for finer meshes. One can also confirm that by enhancing training samples in the case of a decent improvement in overall prediction accuracy is achieved. On the other hand, in Fig. 26 the error started to blow up after several time steps when 3000 training samples were utilized for training in the case of . This was significantly mitigated by enhancing the training samples from 3000 to 5000, part of which is through adding higher frequencies to the Fourier series sample generator. This indicates that in the heterogeneous case, the quality of training samples critically affects the prediction accuracy in the finer mesh as shown in the case of , which is different from the observation in Fig. 19 where one does not see such an extreme error evolution. When it comes to the training cost, one has to pay the price for increasing the number of nodes in FOL. As shown in Fig. 28, the training time increased linearly with increasing number of nodes. The homogeneous case with required 2.645 times more training time than the same setup with . The same trend was confirmed for the heterogeneous case. It is noted that the training was done within 30 minutes for all four cases using JAX. On the other hand, the more the number of nodes increases, the more speedup FOL can achieve against FEM as shown in Fig. 28, given the same network architecture with varying input and output dimensions. This indicates the strong potential of FOL as a fast surrogate to conventional FE solvers for large models with a large number of nodes. In application scenarios, one has to therefore adapt the variety of training samples to achieve reasonable prediction, especially when considering heterogeneous domains. Further exploring the integration of this method with data-driven auto-encoders presents intriguing possibilities to address potential constraints associated with increased mesh densities (i.e. higher resolutions), as highlighted in [75, 78, 79].
@@ 4.7 Capability of handling arbitrary domains
shown in Figs. 30 and 31, the overall solution trend obtained by the proposed FOL framework agreed with the solution by FEM with the maximum absolute error around 0.1. The error was concentrated in the upper right and left part of the domain at and [s], which is also explained by the steep change in the temperature evolution due to the presence of heterogeneity. Furthermore, the error magnitudes did not differ very much from the results with the square domain as in Figs. 13, 14, and 15. The results of Fig. 31 are particularly noteworthy. This proves that FOL works for a problem setup with complex initial temperature, heterogeneity, and irregular domain discretized by unstructured mesh without cumbersome modification to the framework.
@@ 4.8 Computational cost and advantages of proposed framework
The main goal of this work is to establish a surrogate model for conventional numerical solvers. In this context, the prediction by the networks should be faster than the solution by numerical analysis. To quantitatively evaluate the speed of obtaining solutions by the FOL framework, the runtimes of the network inference with the same network architecture and finite element calculation were measured by performing the same task. The measurement was performed on the same CPU platform and environment to ensure the fairness of the measurement. We assumed ten-time steps in solving the heat equation with FEM. Therefore, the same network evaluation was performed ten times as well. As a result, the prediction time with the separated network architecture was 10.8 times faster than FEM for the same setup. This result suggests that the network has the potential to be used as a surrogate for classic numerical solvers.
Although a faster inference than solving with a classical solver can be achieved with the proposed framework, one has to train the NNs for a relatively long time. For example, the training time for 1000 epochs on the problem setup shown in Figs. 4 and 7 took approximately 7 hours on a single GPU node of NVIDIA GeForce RTX 2080 12GB. However, the training time will be much faster using JAX (high-performance machine learning framework) which has superior features for deep learning, such as just-in-time compilation and vectorization. In addition, training NNs in FOL is usually a one-time investment. Once trained, users can use it for any input and can obtain the solution much quicker than numerical solvers, even for a model that requires many nodes to solve accurately.
@@ 5 Conclusion
This study has presented a novel physics-informed operator learning framework based on the finite element discretization scheme for spatiotemporal PDEs. After training with various temperature fields, including those generated by Gaussian distribution and Fourier series, as well as constant temperature fields, the network can accurately predict dynamic temperature evolutions for any arbitrary temperature input within the assumed temperature range. This is achieved with a relative L2 error below 0.1 in most cases, without the need for retraining under fixed boundary conditions and domain. The applicability of the method to heterogeneous heat conductivity and irregular domains is also confirmed. Additionally, the suggested network design can achieve over ten times the speed of the corresponding FEM solver on the same platform. It is important to note that the training is conducted entirely without ground truth data obtained from numerical simulations, making the framework a completely unsupervised learning approach. Furthermore, the training efficiency is enhanced compared to other operator learning approaches that rely on time-consuming automatic differentiation. This is because the current framework uses FE-based discretization for space and backward difference approximation for time. To summarize, this work explores the development of deep learning-based surrogates for dynamic physical phenomena without the need for supervised learning.
On the other hand, although the proposed framework offers useful features, there are still some limitations that could be addressed in future work. Firstly, heat conductivity can also be a target of training in addition to the temperature field. This makes the framework flexible for various micro morphologies with phase-field modeling in mind. It may be possible to improve accuracy by implementing a higher-order temporal discretization scheme. One could also think about different network architectures that take multiple temperature fields as input. Moreover, this study focused on transient heat conduction to showcase the performance of the framework. The present framework could be extended to other types of spatiotemporal PDEs, such as the convection-diffusion equation, the Allen-Cahn equation, or the Cahn-Hilliard equation, as the framework is developed with the aim of a generic deep learning framework for spatiotemporal dynamics phenomena described by PDEs. Additionally, although the present study only utilizes the bilinear interpolation that works for the majority of the possible applications, this framework could also be combined with higher-order basis functions such as the quadratic one for a better representation of geometry with curvature and further accuracy in prediction. Lastly, to efficiently handle large models with a large number of nodes, a reduced parametric space illustrated in Fig. 32 could be introduced by employing techniques such as autoencoder [78, 79, 80].
Author Statement:
A.H. and Sh.R. conceptualized the study. Y.Y., A.H., and Sh.R. developed the methodology. Y.Y. and Sh.R. developed the software. M.M. provided the computational resources for conducting the study. Y.Y. conducted the formal analysis and investigation. Y.Y., A.H., and Sh.R. wrote the original draft. All authors reviewed and edited the manuscript. Sh.R. supervised the project.
Competing Interests:
The authors declare no competing financial or non-financial interests.